# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 2.4.1 Trapped ions
We start with the trapped-ion experiment reported in Ref. Burd et al. (2021); the experimental setup is shown in Fig. 7. The experiment used two ${}^{25}{\rm Mg}^{+}$ ions, which were trapped $\sim 30\,\mathrm{\mu m}$ above a linear surface-electrode ion trap [see Figs. 7(a) and 7(b)]. Furthermore, an out-of-phase radial motional mode, shared by the two ${}^{25}{\rm Mg}^{+}$ ions and cooled to near the ground state using resolved-sideband cooling from oscillating magnetic field gradients, was used as a bosonic harmonic oscillator degree of freedom, and therefore played the role of light in light-matter interactions. Its parametric modulation, and in turn its squeezing, was implemented by applying an oscillating potential at or close to twice the motional frequency to the rf electrodes of the ion trap [see Figs. 7(b) and 7(c)]. The states $\left|{\downarrow}\right\rangle\equiv\left|{F=3,m_{F}=1}\right\rangle$ and $\left|{\uparrow}\right\rangle\equiv\left|{F=2,m_{F}=1}\right\rangle$ in the ${}^{2}S_{1/2}$ electronic ground state hyperfine manifold of the trapped ${}^{25}{\rm Mg}^{+}$ ions were used as qubit states. Here, $F$ refers to the total angular momentum, and $m_{F}$ refers to the projection of the total angular momentum along a quantization axis defined by an external magnetic field.
@@
In the experiment in Ref. Burd et al. (2021), the observed results demonstrated that with parametric amplification, the entanglement generation is sped up by a factor of $\simeq 3.74$, from which an increase of up to $3.25$-fold in the strength of the ion-motion coupling can be estimated. That experiment also showed the dependence of the interaction amplification on the phase $\theta_{\rm 2ph}$, i.e., on the direction of squeezing. This is different from the $\theta_{\rm 2ph}$-independent amplification process proposed in Ref. Qin et al. (2018).
In proposals Lü et al. (2015); Lemonde et al. (2016); Qin et al. (2018); Leroux et al. (2018); Zeytinoğlu et al. (2017) of using parametric squeezing to enhance light-matter interactions, precise knowledge of the system Hamiltonian is needed in advance, e.g., the relative phase between the squeezing operation and the rest of the system dynamics. To clarify this issue, the use of parametric squeezing to amplify a coherent displacement of a bosonic field can be taken as an example. A direct method is to apply a squeezing operation and then an antisqueezing operation to the bosonic field, yielding
@@
where $\phi_{\alpha}={\rm arg}(\alpha)$. Clearly, choosing $\theta_{\rm 2ph}-2\phi_{\alpha}=2n\pi$ ($n=0,\pm 1,\pm 2,\cdots$) can lead to an exponential amplification in the coherent amplitude $\alpha$, such that $\alpha_{\rm sq}=e^{r}\alpha$. This kind of amplification was experimentally reported in Ref. Burd et al. (2019), using the same setup as shown in Fig. 7, but with a single trapped ${}^{25}{\rm Mg}^{+}$ ion. The displacement operation of motion was implemented by applying a resonant oscillating potential to an electrode of the ion trap. An increase in $\alpha$ by a factor of $\approx 9.17$ was demonstrated in that experimental work. However, it is also clear that the amplification given in Eq. (97) depends strongly on the phase difference $\theta_{\rm 2ph}-2\phi_{\alpha}$. Indeed, for some values of $\theta_{\rm 2ph}-2\phi_{\alpha}$, the coherent amplitude $\alpha$ even becomes deamplified; e.g., the phase difference of $\theta_{\rm 2ph}-2\phi_{\alpha}=(2n+1)\pi$ ($n=0,\pm 1,\pm 2,\cdots$) leads to $\alpha_{\rm sq}=e^{-r}\alpha$, i.e., an exponential decrease in $\alpha$. Thus, in order to achieve the desired amplification, the accurate value of the phase $\phi_{\alpha}$ is needed. However, in some cases, the phase $\phi_{\alpha}$ may be unknown or even fluctuate in time. To address this kind of issue, a proposal, which, in the absence of the precise knowledge of the system parameters, can amplify or speed up quantum dynamics using parametric squeezing, was put forward in Ref. Arenz et al. (2020).
Figure 9: Phase-space representation of squeezing-induced amplification of a coherent displacement of motion, demonstrated in Ref. Burd et al. (2023). The gray dotted outlines and the orange blobs represent the phase-space distributions before and after, respectively, the operations shown in the corresponding boxes at the top are applied. The gray and blue arrows refer to the squeezing and displacement operations, respectively, while the red and black arrows in the final panel refer to the unamplified and amplified displacements, respectively. All panels are reproduced with permission from Ref. Burd et al. (2023), S. C. Burd et al., arXiv:2304.05529.
In order to amplify the coherent amplitude $\alpha$ of a bosonic field, whose full knowledge is unknown, the expression in Eq. (96) needs to be modified, according to the formalism of Ref. Arenz et al. (2020), to become
@@
More generally, the method presented in Ref. Arenz et al. (2020) can also amplify a qubit-field interaction without precise knowledge. Consider, as an example, the JC interaction $H_{\rm JC}^{\rm int}$ given in Eq. (54). The evolution operator $U(t)$ under $H_{\rm JC}^{\rm int}$ for an evolution time $t$ is Trotterized, such that
@@
where $U_{0}(t/2N)=\exp(-iH_{\rm JC}^{\rm int}t/2N)$ accounts for the evolution operator under $H_{\rm JC}^{\rm int}$ for a sufficiently small time $t/2N$ Burd et al. (2023); Arenz et al. (2020). According to the Trotter formula Arenz et al. (2020), a sufficiently large $N$ can enable
@@
Recently, this kind of squeezing-induced amplification in the absence of a full knowledge of the details of the system parameters has been experimentally reported in trapped-ion systems in Ref. Burd et al. (2023). The experimental setup was the same as given in Fig. 7, but a single trapped ${}^{25}{\rm Mg}^{+}$ ion was used there. The theoretical predictions given in Eqs. (99) and (103) were demonstrated in that experimental work. The experiment showed that the coherent amplitude $\alpha$ can be amplified by a mean gain of $\approx 1.77$, in good agreement with the theoretical prediction of $\approx 2.11$ (corresponding to a squeezing parameter of $r\approx 1.38$).
@@ 2.4.2 Superconducting circuits
Figure 10: Experimental demonstration of the squeezing-enhanced dispersive atom-cavity coupling reported in Ref. Villiers et al. (2022). (a) Experimental device with a diagonal break. A quarter wavelength coplanar waveguide resonator (blue), as a single-mode cavity, is shunted to ground through a SNAIL element (boxed in green), such that it can act as a degenerate parametric amplifier, whose reflection spectrum is measured through an inductive coupler (boxed in blue). This SNAIL resonator is capacitively coupled to a transmon qubit (red) with a SQUID loop (boxed in red) threaded by a flux. (b) Dispersive coupling $\chi_{\rm trans}$ of the qubit and the cavity versus the two-photon driving $\Omega_{\rm 2ph}$ for $\Delta_{\rm 2ph}/2\pi=0$, ±$\pm 20$, ±$\pm 30$, and ±$\pm 40\,\mathrm{MHz}$, and with offsets $0.2$, $0.1$, and $0\,\mathrm{MHz}$ for clarity. The colormap represents the squeezing in decibels, and the gray shaded area refers to the instability regime. The curves are theoretical predictions, and the symbols are experimental data; moreover, in the case of $\Delta_{\rm 2ph}\neq 0$, the open and solid symbols correspond to $\Delta_{\rm 2ph}>0$ and $\Delta_{\rm 2ph}<0$, respectively. Clearly, there is no improvement in $\chi_{\rm trans}$ for $\Delta_{\rm 2ph}=0$, but a two-fold increase in $\chi_{\rm trans}$ for $\Delta_{\rm 2ph}/2\pi=20\,\mathrm{MHz}$ at a squeezing of $\simeq 5.5\,\mathrm{dB}$. Panels (a) and (b) are reproduced with permission from Ref. Villiers et al. (2022), M. Villiers et al., arXiv:2212.04991.
In the superconducting-circuit experiment in Ref. Villiers et al. (2022), a quarter-wavelength coplanar-waveguide resonator interrupted by a superconducting nonlinear asymmetric inductive element (SNAIL) is capacitively coupled to a transmon qubit, as shown in Fig. 10(a). This SNAIL resonator, when pumped close to twice the resonator frequency, can behave as a degenerate parametric amplifier, and as a result can induce a significant squeezing but with a very weak Kerr nonlinearity.
@@
In the case of $\Delta_{\rm 2ph}=0$, the amplification bandwidth decreases as the gain increases. This is because of the fundamental gain-bandwidth constraint; that is, the product of the gain and the amplification bandwidth is approximately equal to the single-photon loss rate $\kappa$. In sharp contrast, in the detuned case of $\Delta_{\rm 2ph}\neq 0$, the amplification bandwidth becomes constant and thus independent of the gain Metelmann et al. (2022). This interesting feature was also observed in the experiment in Ref. Villiers et al. (2022).
@@ 3 Simulation of ultrastrong and deep-strong light-matter interactions
In this section, we switch from considering schemes for amplifying existing interactions to ways of simulating interactions that are significant compared to the bare transition frequencies in the system. There are many ways to perform such quantum simulations Buluta and Nori (2009); Georgescu et al. (2014), and a plethora of experimental systems in which it can be implemented. Here, we first consider, in Secs. 3.13.3, analog simulation schemes where one or two drives are added to a system with a low interaction strength, such that a reference frame can be found where the renormalized parameters can be tuned by the external drives to reach the USC or DSC regimes. Then, in Sec. 3.4, we review digital protocols for quantum simulation. Finally, in Sec. 3.5, we review a number of additional simulation methods and setups, including VQS, two analog simulation setups with cold atoms, and ways to simulate the USC between resonators or in optomechanical systems.
@@ 3.1.1 Derivation
Some of the most important instances of simulating light-matter systems in the USC regime are based on the use of cavity-assisted Raman transitions to obtain an effective Dicke Hamiltonian. Crucially, the effective frequencies of the involved subsystems in the simulated Hamiltonian are strongly reduced, thereby providing a way to explore regimes where the light-matter coupling can be very strong in comparison. Such a simulation was originally proposed in Ref. Dimer et al. (2007) as a way to observe the superradiant phase transition of the Dicke model, and has been followed by many proposals for particular systems. For example, in Ref. Grimsmo and Parkins (2013), a method was proposed to simulate the quantum Rabi model in cavity QED with a single rubidium atom in a cavity.
Figure 11: Level structures of simulations of the quantum Rabi model with cavity-assisted Raman transitions. (a) Original proposal with the atomic electronic states, and (b) alternative proposal with the atomic motional states in a BEC. See also Refs. Dimer et al. (2007) and Baumann et al. (2010), respectively.
@@
Under the assumption $\left|\Delta_{r/s}\right|\gg\{\delta_{c},\Delta_{1},\Omega_{r/s},g_{r/s}\}$, the excited states can then be adiabatically eliminated. This assumption enables a separation between a manifold of “fast” degrees of freedom (involving the excited states), and a manifold of “slow” degrees of freedom (involving the ground states). After adiabatically eliminating the fast degrees of freedom, an effective Hamiltonian for the slow degrees of freedom can be obtained: the states $\left|{1}\right\rangle$ and $\left|{0}\right\rangle$ are connected via the second-order processes, which involve the absorption-emission processes of a cavity photon and a laser photon. As an intuitive example, one can see that the excitation $\left|{0}\right\rangle\rightarrow\left|{r}\right\rangle$ by the absorption of a cavity photon, followed by the transition $\left|{r}\right\rangle\rightarrow\left|{1}\right\rangle$ via the emission of a laser photon, yields an effective dynamics described by a Hamiltonian term $\propto a\sigma_{01}^{+}$. In order to derive the effective Hamiltonian, one can follow the approach based on the Schrieffer-Wolff transformation described in Ref. Cohen‐Tannoudji et al. (1998). Up to second order in the interaction term $V$ that couples fast and slow degrees of freedom, the effective Hamiltonian is expressed as
@@
While we here have presented the derivation for the one-atom case, the effective Hamiltonian for a system of $N$ atoms is only slightly modified by including a factor $N$ for the Lamb shift of the cavity resonance, i.e., $\lambda_{A}\rightarrow N\lambda_{A}$ Dimer et al. (2007). It can be seen that one can reach $\chi=0$ by setting $g_{r}^{2}/\Delta_{r}=g_{s}^{2}/\Delta_{s}$. Similarly, we can set $g_{s}\Omega_{s}/\Delta_{s}=g_{r}\Omega_{r}/2\Delta_{r}$, so that the effective Hamiltonian has the form of a quantum Rabi model for $N=1$, and a quantum Dicke model for $N>1$:
@@ 3.1.2 Implementations in atomic condensates
The first experimental implementation of cavity-assisted Raman transitions for the simulation of the Dicke model, and also the observation of a superradiant transition, was realized in an atomic BEC Baumann et al. (2010). In this approach, the spin is defined by the atomic motional states, rather than the atomic electronic states. The underlying mechanism behind the observation of this superradiant transition in atomic clouds inside optical cavities is the self-organization of the atoms into a checkerboard pattern, as demonstrated theoretically in Ref. Domokos and Ritsch (2002) and shown experimentally in Ref. Black et al. (2003). The theoretical mapping of this transition to a dynamical version of the superradiance transition of the Dicke model was put forward in Refs. Chen et al. (2007); Nagy et al. (2010). The equivalence between them is established by writing the state of an atom as $\left|{k_{x},k_{z}}\right\rangle\left|{a}\right\rangle$, where $\left|{k_{x},k_{z}}\right\rangle$ describes a momentum eigenstate in the $x$-$z$ plane and $\left|{a}\right\rangle$ is an internal, electronic eigenstate. The energy-level structure and the corresponding couplings are shown in Fig. 11(b).
@@
This idea was implemented experimentally in Ref. Baumann et al. (2010), which reported the first Dicke phase transition, using a ${}^{87}\mathrm{Rb}$ BEC with $N\sim 10^{5}$ atoms. The Dicke phase transition and the emergence of self-organization are manifested by an abrupt build-up of the cavity field accompanied by the development of momentum components at $(k_{x},k_{z})$ in the atomic cloud. The symmetry breaking between two possible superradiant states (associated with two atomic density waves shifted by half an optical wavelength) was further explored by the same group in Ref. Baumann et al. (2011), and the dynamical properties of this phase transition were then interpreted in terms of a Kibble-Zurek mechanism in Ref. Klinder et al. (2015). Several theoretical works have studied more deeply the critical properties and phase diagrams of these models; see, e.g., Refs. Bhaseen et al. (2012); Kirton and Keeling (2018).

### v2

@@ 2.4.1 Trapped ions
We start with the trapped-ion experiment reported in Ref. [206]; the experimental setup is shown in Fig. 7. The experiment used two ${}^{25}{\rm Mg}^{+}$ ions, which were trapped $\sim 30\,\mathrm{\mu m}$ above a linear surface-electrode ion trap [see Figs. 7(a) and 7(b)]. Furthermore, an out-of-phase radial motional mode, shared by the two ${}^{25}{\rm Mg}^{+}$ ions and cooled to near the ground state using resolved-sideband cooling from oscillating magnetic field gradients, was used as a bosonic harmonic oscillator degree of freedom, and therefore played the role of light in light-matter interactions. Its parametric modulation, and in turn its squeezing, was implemented by applying an oscillating potential at or close to twice the motional frequency to the rf electrodes of the ion trap [see Figs. 7(b) and 7(c)]. The states $\left|{\downarrow}\right\rangle\equiv\left|{F=3,m_{F}=1}\right\rangle$ and $\left|{\uparrow}\right\rangle\equiv\left|{F=2,m_{F}=1}\right\rangle$ in the ${}^{2}S_{1/2}$ electronic ground state hyperfine manifold of the trapped ${}^{25}{\rm Mg}^{+}$ ions were used as qubit states. Here, $F$ refers to the total angular momentum, and $m_{F}$ refers to the projection of the total angular momentum along a quantization axis defined by an external magnetic field.
@@
In the experiment in Ref. [206], the observed results demonstrated that with parametric amplification, the entanglement generation is sped up by a factor of $\simeq 3.74$, from which an increase of up to $3.25$-fold in the strength of the ion-motion coupling can be estimated. That experiment also showed the dependence of the interaction amplification on the phase $\theta_{\rm 2ph}$, i.e., on the direction of squeezing. This is different from the $\theta_{\rm 2ph}$-independent amplification process proposed in Ref. [289].
In proposals [246, 256, 289, 290, 291] of using parametric squeezing to enhance light-matter interactions, precise knowledge of the system Hamiltonian is needed in advance, e.g., the relative phase between the squeezing operation and the rest of the system dynamics. To clarify this issue, the use of parametric squeezing to amplify a coherent displacement of a bosonic field can be taken as an example. A direct method is to apply a squeezing operation and then an antisqueezing operation to the bosonic field, yielding
@@
where $\phi_{\alpha}={\rm arg}(\alpha)$. Clearly, choosing $\theta_{\rm 2ph}-2\phi_{\alpha}=2n\pi$ ($n=0,\pm 1,\pm 2,\cdots$) can lead to an exponential amplification in the coherent amplitude $\alpha$, such that $\alpha_{\rm sq}=e^{r}\alpha$. This kind of amplification was experimentally reported in Ref. [208], using the same setup as shown in Fig. 7, but with a single trapped ${}^{25}{\rm Mg}^{+}$ ion. The displacement operation of motion was implemented by applying a resonant oscillating potential to an electrode of the ion trap. An increase in $\alpha$ by a factor of $\approx 9.17$ was demonstrated in that experimental work. However, it is also clear that the amplification given in Eq. (97) depends strongly on the phase difference $\theta_{\rm 2ph}-2\phi_{\alpha}$. Indeed, for some values of $\theta_{\rm 2ph}-2\phi_{\alpha}$, the coherent amplitude $\alpha$ even becomes deamplified; e.g., the phase difference of $\theta_{\rm 2ph}-2\phi_{\alpha}=(2n+1)\pi$ ($n=0,\pm 1,\pm 2,\cdots$) leads to $\alpha_{\rm sq}=e^{-r}\alpha$, i.e., an exponential decrease in $\alpha$. Thus, in order to achieve the desired amplification, the accurate value of the phase $\phi_{\alpha}$ is needed. However, in some cases, the phase $\phi_{\alpha}$ may be unknown or even fluctuate in time. To address this kind of issue, a proposal, which, in the absence of the precise knowledge of the system parameters, can amplify or speed up quantum dynamics using parametric squeezing, was put forward in Ref. [260].
Figure 9: Phase-space representation of squeezing-induced amplification of a coherent displacement of motion, demonstrated in Ref. [207]. The gray dotted outlines and the orange blobs represent the phase-space distributions before and after, respectively, the operations shown in the corresponding boxes at the top are applied. The gray and blue arrows refer to the squeezing and displacement operations, respectively, while the red and black arrows in the final panel refer to the unamplified and amplified displacements, respectively. All panels are reproduced from Ref. [207], S. C. Burd et al., PRX Quantum 5, 020314 (2024), with permission.
In order to amplify the coherent amplitude $\alpha$ of a bosonic field, whose full knowledge is unknown, the expression in Eq. (96) needs to be modified, according to the formalism of Ref. [260], to become
@@
More generally, the method presented in Ref. [260] can also amplify a qubit-field interaction without precise knowledge. Consider, as an example, the JC interaction $H_{\rm JC}^{\rm int}$ given in Eq. (54). The evolution operator $U(t)$ under $H_{\rm JC}^{\rm int}$ for an evolution time $t$ is Trotterized, such that
@@
where $U_{0}(t/2N)=\exp(-iH_{\rm JC}^{\rm int}t/2N)$ accounts for the evolution operator under $H_{\rm JC}^{\rm int}$ for a sufficiently small time $t/2N$ [207, 260]. According to the Trotter formula [260], a sufficiently large $N$ can enable
@@
Recently, this kind of squeezing-induced amplification in the absence of a full knowledge of the details of the system parameters has been experimentally reported in trapped-ion systems in Ref. [207]. The experimental setup was the same as given in Fig. 7, but a single trapped ${}^{25}{\rm Mg}^{+}$ ion was used there. The theoretical predictions given in Eqs. (99) and (103) were demonstrated in that experimental work. The experiment showed that the coherent amplitude $\alpha$ can be amplified by a mean gain of $\approx 1.77$, in good agreement with the theoretical prediction of $\approx 2.11$ (corresponding to a squeezing parameter of $r\approx 1.38$).
@@ 2.4.2 Superconducting circuits
Figure 10: Experimental demonstration of the squeezing-enhanced dispersive atom-cavity coupling reported in Ref. [209]. (a) Experimental setup with a diagonal break. A quarter wavelength coplanar waveguide resonator (blue), as a single-mode cavity, is shunted to ground through a SNAIL element (boxed in green), such that it can act as a degenerate parametric amplifier, whose reflection spectrum is measured through an inductive coupler (boxed in blue). This SNAIL resonator is capacitively coupled to a transmon qubit (red) with a SQUID loop (boxed in red) threaded by a flux. (b) Dispersive coupling $\chi_{\rm trans}$ of the qubit and the cavity versus the two-photon driving $\Omega_{\rm 2ph}$ for $\Delta_{\rm 2ph}/2\pi=0$, ±$\pm 20$, ±$\pm 30$, and ±$\pm 40\,\mathrm{MHz}$, and with offsets $0.2$, $0.1$, and $0\,\mathrm{MHz}$ for clarity. The colormap represents the squeezing in decibels, and the gray shaded area refers to the instability regime. The curves are theoretical predictions, and the symbols are experimental data; moreover, in the case of $\Delta_{\rm 2ph}\neq 0$, the open and solid symbols correspond to $\Delta_{\rm 2ph}>0$ and $\Delta_{\rm 2ph}<0$, respectively. Clearly, there is no improvement in $\chi_{\rm trans}$ for $\Delta_{\rm 2ph}=0$, but a two-fold increase in $\chi_{\rm trans}$ for $\Delta_{\rm 2ph}/2\pi=20\,\mathrm{MHz}$ at a squeezing of $\simeq 5.5\,\mathrm{dB}$. Figures, with some modifications, from Ref. [209], M. Villiers et al., PRX Quantum 5, 020306 (2024), with permission.
In the superconducting-circuit experiment in Ref. [209], a quarter-wavelength coplanar-waveguide resonator interrupted by a superconducting nonlinear asymmetric inductive element (SNAIL) is capacitively coupled to a transmon qubit, as shown in Fig. 10(a). This SNAIL resonator, when pumped close to twice the resonator frequency, can behave as a degenerate parametric amplifier, and as a result can induce a significant squeezing but with a very weak Kerr nonlinearity.
@@
In the case of $\Delta_{\rm 2ph}=0$, the amplification bandwidth decreases as the gain increases. This is because of the fundamental gain-bandwidth constraint; that is, the product of the gain and the amplification bandwidth is approximately equal to the single-photon loss rate $\kappa$. In sharp contrast, in the detuned case of $\Delta_{\rm 2ph}\neq 0$, the amplification bandwidth becomes constant and thus independent of the gain [369]. This interesting feature was also observed in the experiment in Ref. [209].
@@ 3 Simulation of ultrastrong and deep-strong light-matter interactions
In this section, we switch from considering schemes for amplifying existing interactions to ways of simulating interactions that are significant compared to the bare transition frequencies in the system. There are many ways to perform such quantum simulations [177, 178], and a plethora of experimental systems in which it can be implemented. Here, we first consider, in Secs. 3.13.3, analog simulation schemes where one or two drives are added to a system with a low interaction strength, such that a reference frame can be found where the renormalized parameters can be tuned by the external drives to reach the USC or DSC regimes. Then, in Sec. 3.4, we review digital protocols for quantum simulation. Finally, in Sec. 3.5, we review a number of additional simulation methods and setups, including VQS, two analog simulation setups with cold atoms, and ways to simulate the USC between resonators or in optomechanical systems.
@@ 3.1.1 Derivation
Some of the most important instances of simulating light-matter systems in the USC regime are based on the use of cavity-assisted Raman transitions to obtain an effective Dicke Hamiltonian. Crucially, the effective frequencies of the involved subsystems in the simulated Hamiltonian are strongly reduced, thereby providing a way to explore regimes where the light-matter coupling can be very strong in comparison. Such a simulation was originally proposed in Ref. [370] as a way to observe the superradiant phase transition of the Dicke model, and has been followed by many proposals for particular systems. For example, in Ref. [371], a method was proposed to simulate the quantum Rabi model in cavity QED with a single rubidium atom in a cavity.
Figure 11: Level structures of simulations of the quantum Rabi model with cavity-assisted Raman transitions. (a) Original proposal with the atomic electronic states, and (b) alternative proposal with the atomic motional states in a BEC. See also Refs. [370] and [372], respectively.
@@
Under the assumption $\left|\Delta_{r/s}\right|\gg\{\delta_{c},\Delta_{1},\Omega_{r/s},g_{r/s}\}$, the excited states can then be adiabatically eliminated. This assumption enables a separation between a manifold of “fast” degrees of freedom (involving the excited states), and a manifold of “slow” degrees of freedom (involving the ground states). After adiabatically eliminating the fast degrees of freedom, an effective Hamiltonian for the slow degrees of freedom can be obtained: the states $\left|{1}\right\rangle$ and $\left|{0}\right\rangle$ are connected via the second-order processes, which involve the absorption-emission processes of a cavity photon and a laser photon. As an intuitive example, one can see that the excitation $\left|{0}\right\rangle\rightarrow\left|{r}\right\rangle$ by the absorption of a cavity photon, followed by the transition $\left|{r}\right\rangle\rightarrow\left|{1}\right\rangle$ via the emission of a laser photon, yields an effective dynamics described by a Hamiltonian term $\propto a\sigma_{01}^{+}$. In order to derive the effective Hamiltonian, one can follow the approach based on the Schrieffer-Wolff transformation described in Ref. [373]. Up to second order in the interaction term $V$ that couples fast and slow degrees of freedom, the effective Hamiltonian is expressed as
@@
While we here have presented the derivation for the one-atom case, the effective Hamiltonian for a system of $N$ atoms is only slightly modified by including a factor $N$ for the Lamb shift of the cavity resonance, i.e., $\lambda_{A}\rightarrow N\lambda_{A}$ [370]. It can be seen that one can reach $\chi=0$ by setting $g_{r}^{2}/\Delta_{r}=g_{s}^{2}/\Delta_{s}$. Similarly, we can set $g_{s}\Omega_{s}/\Delta_{s}=g_{r}\Omega_{r}/2\Delta_{r}$, so that the effective Hamiltonian has the form of a quantum Rabi model for $N=1$, and a quantum Dicke model for $N>1$:
@@ 3.1.2 Implementations in atomic condensates
The first experimental implementation of cavity-assisted Raman transitions for the simulation of the Dicke model, and also the observation of a superradiant transition, was realized in an atomic BEC [372]. In this approach, the spin is defined by the atomic motional states, rather than the atomic electronic states. The underlying mechanism behind the observation of this superradiant transition in atomic clouds inside optical cavities is the self-organization of the atoms into a checkerboard pattern, as demonstrated theoretically in Ref. [374] and shown experimentally in Ref. [375]. The theoretical mapping of this transition to a dynamical version of the superradiance transition of the Dicke model was put forward in Refs. [376, 377]. The equivalence between them is established by writing the state of an atom as $\left|{k_{x},k_{z}}\right\rangle\left|{a}\right\rangle$, where $\left|{k_{x},k_{z}}\right\rangle$ describes a momentum eigenstate in the $x$-$z$ plane and $\left|{a}\right\rangle$ is an internal, electronic eigenstate. The energy-level structure and the corresponding couplings are shown in Fig. 11(b).
@@
This idea was implemented experimentally in Ref. [372], which reported the first Dicke phase transition, using a ${}^{87}\mathrm{Rb}$ BEC with $N\sim 10^{5}$ atoms. The Dicke phase transition and the emergence of self-organization are manifested by an abrupt build-up of the cavity field accompanied by the development of momentum components at $(k_{x},k_{z})$ in the atomic cloud. The symmetry breaking between two possible superradiant states (associated with two atomic density waves shifted by half an optical wavelength) was further explored by the same group in Ref. [378], and the dynamical properties of this phase transition were then interpreted in terms of a Kibble-Zurek mechanism in Ref. [379]. Several theoretical works have studied more deeply the critical properties and phase diagrams of these models; see, e.g., Refs. [380, 381].